OpenStudy (anonymous):

help me find this please i know its simple, i am just having a hard time understanding what it is asking! let the integral (-2,7) fxdx=10, integral (-2,1) fxdx=7, integral (4,7) fxdx=3, Find the integral (1,4) (10fx-7)dx

OpenStudy (amistre64):

(-2,7) - (-2,1) = (1,7); (1,7)-(4,7) = (1,4)

OpenStudy (amistre64):

10* (1,4)fx dx- (1,4) 7dx

OpenStudy (amistre64):

10 - 7 = 3; 3 - 3 = 0 10*(1,4)fx dx = 0 -7x; (1,4) = -7(4) - -7(1) = -28 + 7 = -21

OpenStudy (amistre64):